# PePNet: A Periodicity-Perceived Workload Prediction Network Supporting Rare Occurrence of Heavy Workload

### v2

@@ III-B Prediction Accuracy
We summarize the performance of all the methods in Tab. IV. We use the above metrics to evaluate the accuracy of heavy-workload forecasting accuracy as well as overall forecasting accuracy. We highlight the highest accuracy in boldface. Besides, if PePNet achieves the highest accuracy, we highlight the best one in baseline with underlines.
As shown in Tab. IV, PePNet achieves the best heavy-workload prediction accuracy as well as overall prediction accuracy on all three workload datasets, with few exceptions. Comparing the prediction accuracy on all three datasets, PePNet works the best on long-tail distributed Dinda’s dataset, while less advantageous on the relatively even-distributed Alibaba2018-CPU dataset.
Long-tail data distribution. PePNet works best in this case. Compared with the best performance among the baselines, PePNet improves MAPE, MSE, and MAE by 17.9%, 24.0%, and 29.5% respectively for overall prediction. For heavy-workload prediction, PePNet improves MAPE, MSE, and MAE by 40.8%, 10.4%, and 38.7% respectively over the best method in the baseline.
Centralized data distribution with rare extreme value. PePNet also has obvious advantages in this case, because the SMD dataset has extremely heavy workload. Compared with the best performance in the baseline, PePNet improves MAPE, MSE and MAE by 14.0%, 25.2% and 26.7% respectively for overall prediction. For heavy-workload prediction, PePNet improves MAPE and MAE by 25.3% and 25.0% respectively.
Centralized data distribution. PePNet is also effective in this case. Compared with the best performance in the baseline, PePNet improves MAPE, MSE and MAE by 33.8%, 13.9% and 9.5% for overall prediction. For heavy-workload prediction, PePNet improves MAPE, MSE and MAE by 15.2%, 24.4% and 8.7% respectively, compared with the best method in the baseline.
Relatively even data distribution. In this case, PePNet maintains high overall prediction accuracy and achieves the best heavy workload prediction performance among neural network-based methods. PePNet improves MAPE, MSE and MAE by 25.7%, 28.3% and 13.9% for heavy-workload prediction over the best method in the baseline.
@@ III-D Hyperparameter Sensitivity
The impact of input length and prediction length. We use grid search to explore the impact of $X^{enc}_{short}$’s length and the impact of $y$’s length on PePNet’s performance. We show the result in Fig. 4(d)-Fig. 4(i). We perform experiments with Cartesian combinations of input lengths from 30 to 80 and prediction lengths from 2 to 12. Generally, the error grows higher gently when predicting length increases. The best input length for different datasets is different. It is found the best input length for Alibaba2018-Memory and Alibaba2018-CPU is 50, while the best input length for Dinda’s dataset is 30.
@@
Dinda’s dataset. For overall workload prediction, MAE rises slowly as prediction length increases and MAE is stable for any input length. For heavy-workload workload prediction, MAE is more stable for prediction length when the input length is longer. As the length of prediction grows, the MAE for overall accuracy increases no more than 0.051 at each input length, while the MAE of heavy workload increases no more than 0.062.
@@
Alibaba2018-Memory. The long-term dependency of Alibaba’s memory usage is weak, thus when the input length is too big, increasing the input length increases the noise level, which could corrupt the performance of PePNet. Thus, when the input length is greater than 50, the MAE of PePNet is unstable. But when the input length is less than 50, the performance of PePNet is stable, and the prediction error only increases slowly with the prediction length. For overall memory usage prediction, the MAE with a prediction length of 12 is only 0.006 larger than the MAE with a prediction length of 2. For heavy-workload memory usage prediction, the MAE with a prediction length of 12 is only 0.012 larger than that with a prediction length of 2.
@@
Alibaba2018-CPU. Long-term dependency is more pronounced in CPU usage. Thus, when we increase the input length, the performance of PePNet is always stable. For overall CPU usage prediction, the MAE with a prediction length of 12 is only 0.016 larger than that with a prediction length of 2. For heavy-workload CPU usage prediction, the MAE with a prediction length of 12 is only 0.034 larger than that with a prediction length of 2.
@@
The impact of $\gamma$. We also explore the impact of $\gamma$ on three datasets, as shown in Fig. 4(c). There are slight fluctuations of overall accuracy for different $\gamma$. Besides, it is consistent with the theoretical analysis in Section.II-C that when the $\gamma$ is smaller the heavy-workload accuracy is higher. It is also intuitive that the heavy-workload MAE of Dinda’s dataset dips most sharply, as Dinda’s dataset is long-tail distributed. The extremely heavy workload in the long tail greatly reduces the prediction accuracy. But as $\gamma$ gets smaller, more attention is put on heavy workload, during the process of training. According to our experimental results, setting $\gamma$ as the value indefinitely approaching zero can lead to the best accuracy. As analyzed in Section.II-C, when $\gamma$ indefinitely approaches zero, Achilles’ loss function only optimizes the part with the highest prediction error. In this situation, PePNet also achieves high prediction accuracy without overall accuracy degradation, which verifies that the Achilles’ loss function helps to avoid overly emphasizing sporadic patterns.
@@ III-E Ablation Experiment
We validate the effect of the Periodicity-Perceived Mechanism and heavy-workload-focused loss function by comparing the performance of PePNet with that of PePNet- and PePNet. We show the performance of these models in Tab. IV and summarize the improvement ratio of PePNet compared with PePNet- and PePNet in Tab. V and Tab. VI respectively. Overall, the Periodicity-Perceived Mechanism contributes more to the prediction accuracy than the Achilles’ Heel loss function. But Achilles’ Heel loss function can improve the accuracy of periodic and aperiodic time series, while the Periodicity-Perceived Mechanism can only promote the accuracy of periodic and lax-periodic data. Furthermore, we also test the prediction accuracy of PePNet on aperiodic data.
The performance of Periodicity-Perceived Mechanism.
Dinda’s dataset and SMD dataset exhibit stricter periodicity than Alibaba2018 does [10]. Thus, the Periodicity-Perceived Mechanism promotes the prediction accuracy most on Dinda’s dataset and SMD dataset. For datasets with lax periodicities, such as memory usage of Alibaba2018, the Periodicity-Perceived Mechanism also promotes overall and heavy-workload prediction accuracy. As for CPU usage of Alibaba2018, which has no significant periodicity on light load but has more significant periodicity on heavy workload, the Periodicity-Perceived Mechanism can promote the heavy-workload prediction accuracy, while the performance of PePNet is about as same as PePNet- on light load. This observation confirms that the Periodicity-Perceived Mechanism has little negative impact on aperiodic data.
TABLE V: Improvement (%) of PePNet over PePNet-.
Overall Heavy-workload
MAPE MSE MAE MAPE MSE MAE
Dinda’s dataset 14.8 36.0 37.1 49.9 -8.7 46.6
SMD dataset 25.7 44.1 30.8 18.7 38.5 21.6
Alibaba2018-Memory 67.6 10.0 9.07 2.53 10.1 2.69
Alibaba2018-CPU -2.1 0.60 0.30 19.4 29.9 15.1
Max. 67.6 44.1 37.1 49.9 38.5 46.6
Min. -2.1 0.59 0.29 2.53 -8.7 2.69
Avg. 26.6 22.7 19.3 22.6 17.5 21.5
The performance of Achilles’ Heel loss function. Our loss function can improve both the overall prediction accuracy and heavy-workload prediction accuracy on all three datasets, compared with PePNet. Except MAPE for heavy-workload prediction accuracy on memory usage of Alibaba2018, all of the evaluation metrics of PePNet are better than that of PePNet. As the Achilles’ Heel Loss function apparently reduces the error for extremely heavy workload in the long tail, which is also proven in Fig. 4(c), the Achilles’ Heel Loss Function improves the accuracy of Dinda’s dataset and SMD dataset more that of Alibaba2018.
TABLE VI: Improvement (%) of PePNet over PePNet.
Overall Heavy-workload
MAPE MSE MAE MAPE MSE MAE
Dinda’s dataset 12.2 39.8 40.4 59.4 16.4 56.6
SMD dataset 19.1 40.9 26.3 12.6 31.0 13.2
Alibaba2018-Memory 39.9 5.28 4.96 -1.3 6.09 0.66
Alibaba2018-CPU 10.3 1.10 2.10 18.9 28.7 13.6
Max. 39.9 40.9 40.4 59.4 31.0 56.6
Min. 10.3 1.12 2.13 -1.3 6.09 0.66
Avg. 20.4 21.8 18.5 22.4 20.6 21.0
TABLE VII: Compare the performance of PePNet on aperiodic and periodic data. The MSE are multiplied by 1000. The MAE are multiplied by 100.
@@
The performance of PePNet on aperiodic data. There is a major concern about whether PePNet still works well on aperiodic data. To test the prediction accuracy on aperiodic data, we divide the Alibaba2018-CPU and Alibaba2018-Memory by periodicity and collect the prediction accuracy for periodic data and aperiodic data respectively. In Tab. VII, there is prediction accuracy for periodic data and aperiodic data. For convenience, the overall accuracy of both periodic data and aperiodic data is also listed at the bottom. On the whole, the accuracy of periodic data is slightly higher than the overall accuracy, while the accuracy of aperiodic data is slightly lower than the overall. However, there is a strange phenomenon the MAPE of periodic data in Alibaba2018-Memory is bigger than the overall accuracy, while its MAE and MSE are much lower than the overall accuracy. That is because, in the Alibaba2018-Memory dataset, the periodic workload is much smaller than the aperiodic one. Thus, even slight errors in periodic data prediction tend to become much larger after the division in the computation of MAPE.
@@ IV Related Work
We first review existing methods for workload and time series prediction. Since PePNet uses periodic information, we also summarize related work for this topic.
IV-A Workload and Time Series Prediction
Time series prediction aims to predict future time series based on the observed workload data, while workload prediction aims to predict future workload based on observed ones. Present research predicts many different workloads, such as the GPU workload [27], CPU utilization [28], virtual desktop infrastructure pool workload [29], disk states [30], database query arrival rates [31] and so on. Researchers make workload predictions for different aims, such as cloud resource allocation [32], autoscaling [28], etc. Additionally, some works proposed an integrated framework of prediction models and other functional modules. For example, Uncertainty-Aware Heuristic Search (UAHS) [33] combines workload prediction models and Bayesian optimization to model the uncertainty and guide virtual machines (VMs) provision strategy. SimpleTS [34] deploys a framework that automatically selects a proper prediction model for a specific time series. However, these methods only focus on improving overall accuracy while neglecting high load accuracy.
As workload prediction is a branch of time series prediction, it is necessary to review the development of time series prediction. In this domain, several frequently discussed issues garner significant attention. One stream is to improve the prediction accuracy of dynamic real-world time series, especially the accuracy of long input sequences. The solutions for this problem can be roughly divided into four categories: the statistic-based methods [7] [35] [36], the recurrent neural networks [37] [38] [39], the transformer-based methods [23] [24] [11] and the integrated methods [10] [40] [5]. The statistic-based methods are efficient but not robust for highly dynamic real-world time series. The recurrent neural networks improve this aspect but suffer from catastrophic forgetting effects. To conquer this drawback, the integrated methods combine attention mechanism [10] and auto-encoder [14] into the recurrent neural network.  Transformer-based methods make further progress by utilizing transformers. Another time series prediction issue that receives widespread attention is the spatiotemporal forecasting problem. In this field, researchers improve the prediction accuracy by leveraging the spatial and temporal information with GNN, CNN, plugin modules and other neural networks [15] [41] [42] [43]. But these methods manually define the spatio-temporal blocks which capture the temporal and spatial correlations. To overcome this limitation, AutoCTS [44] automatically selects competitive heterogeneous spatiotemporal blocks. But these methods also ignore the under-fitting problem of extreme values. Another research direction is improving the robustness of deep networks dealing with time series. The robustness includes five implications: the robustness for missing data, the robustness for the malicious or limited labels, the generalization of the trained model, the robustness toward noise in time series, and the robustness for the irregular data sampling period. For example, Belkhouja et al. propose a Robust Training for Time-Series (ROTS) to improve deep neural network robustness for generalization and noise [45]; Tan et al. propose Dual-Attention Time-Aware Gated Recurrent Unit (DATA-GRU) to manipulate time series sampled in an irregular period [46]; Zhang et al. propose Multivariate Time Series Classification with Attentional Prototypical Network (TapNet) to deal with time series with limited labels [47]; Tang et al. propose network modeling local and global temporal dynamics (LGnet) to deal with a data-missing problem in time series [48]; Luo et al. propose $E^{2}GAN$, which uses adversarial learning to deal with data missing problem [49]; Yao et al. propose Spatial-Temporal Dynamic Network (STDN) to deal with lax-strict periodicity caused by noise [16]; Luo et al. propose Uncertainty-Aware Heuristic Search (UAHS) to deal with uncertain prediction error caused by noise [33]. However, none of these methods focus on dealing with the prediction accuracy of heavy workload.
IV-B Periodicity Information Extraction
In the field of traffic forecasting, there is also significant periodicity and we summarize some recent popular methods below. There are several mechanisms to fuse periodic information. For example, Guo et al. propose a novel attention-based spatial-temporal graph convolution network [13]; Lv et al. use a fully connected layer and take advantage of both CNN and RNN to deal with periodic time series [15]; Chen et al. propose Hop Res-RGNN to deal with periodic patterns [14]. However, these methods do not consider the lax periodicity of time series. Yao et al. propose an attention mechanism to tackle periodic shift [16]. These studies use the priori knowledge of the daily and weekly periodicity of traffic loads. However, in the scenario of service load prediction, there is usually no priori knowledge about the periodicity. Besides, the periodicity of the time series in these studies is fixed, while the periodicities of workload for different machines are variable.
@@ V Conclusion
In this paper, we study the problem of improving overall workload prediction accuracy as well as heavy-workload prediction accuracy and propose PePNet. PePNet makes use of short-term dependent information, long-term tendency information and periodic information. Within PePNet, we propose two mechanisms for better prediction accuracy of workloads: (1) a Periodicity-Perceived Mechanism to guide heavy-workload prediction, which can mine the periodic information and adaptively fuse periodic information for periodic, lax periodic and aperiodic time series; and (2) an Achilles’ Heel Loss Function to offset the negative effect of data imbalance. We also provide theoretical support for the above design by: (1) providing a theoretically proven error bound of periodic information extracted by Periodicity-Perceived Mechanism; and (2) providing an automatic hyperparameter determination method for periodicity threshold $\mathcal{T}$. Compared with existing methods, extensive experiments conducted on Alibaba2018, Dinda’s dataset and SMD dataset demonstrate that PePNet improves MSE for overall workload prediction by 11.8% on average. Especially, PePNet improves MSE for heavy workload prediction by 21.0% on average.
Acknowledgments
This work was supported by the National Science Foundation of China under Grants 62125206.

### v3

@@ III-B Prediction Accuracy
We summarize the performance of all the methods in Tab. II. We highlight the highest accuracy in boldface. Besides, if PePNet achieves the highest accuracy, we highlight the best one among its peers with underlines.
As shown in Tab. II, PePNet achieves the best heavy-workload prediction accuracy as well as overall prediction accuracy on all three workload datasets, with few exceptions. Comparing the prediction accuracy on all three datasets, PePNet works the best on long-tail distributed Dinda’s dataset, while less advantageous on the relatively even-distributed Alibaba2018-CPU dataset.
@@ III-D Hyperparameter Sensitivity
Impact of input length and prediction length. We use grid search [40] to explore the impact of $\bar{X}$’s length and the impact of $y$’s length on PePNet’s performance. We show the result in Fig. 4(d). We perform experiments with Cartesian combinations of input lengths from 30 to 80 and prediction lengths from 2 to 12. Generally, the error grows higher gently when predicting length increases. The best input length for different datasets is different. It is found the best input length for Alibaba2018-Memory and Alibaba2018-CPU is 50, while the best one for Dinda’s dataset is 30.
@@
Dinda’s dataset. For overall workload prediction, MAE rises slowly as prediction length increases and MAE is stable for any input length. For heavy-workload workload prediction, MAE is more stable for prediction length when the input length is longer. As the length of prediction grows, MAE for overall accuracy increases no more than 0.051 at each input length, while MAE of heavy workload prediction increases no more than 0.062.
@@
Alibaba2018-Memory. The long-term dependency of Alibaba’s memory usage is weak, thus when the input length is too big, increasing it increases the noise level, which could corrupt the performance of PePNet. Thus, when the input length is greater than 50, PePNet’s MAE is unstable. But when the input length is less than 50, its performance is stable, and the prediction error only increases slowly with the prediction length. For overall memory usage prediction, MAE with a prediction length of 12 is only 0.006 larger than MAE with a prediction length of 2. For heavy-workload memory usage prediction, MAE with a prediction length of 12 is only 0.012 larger than that with a prediction length of 2.
@@
Alibaba2018-CPU. Long-term dependency is more pronounced in CPU usage. Thus, when we increase the input length, the performance of PePNet is always stable. For overall CPU usage prediction, MAE with a prediction length of 12 is only 0.016 larger than that with a prediction length of 2. For heavy-workload CPU usage prediction, MAE with a prediction length of 12 is only 0.034 larger than that with a prediction length of 2.
@@
Impact of $\gamma$. We explore the impact of $\gamma$ on three datasets, as shown in Fig. 4(c). There are slight fluctuations of overall accuracy for different $\gamma$. Besides, it is consistent with our prior theoretical analysis that when $\gamma$ is smaller the heavy-workload accuracy is higher. It is also intuitive that the heavy-workload MAE of Dinda’s dataset dips most sharply, as Dinda’s dataset is long-tail distributed. The extremely heavy workload in the long tail greatly reduces the prediction accuracy. But as $\gamma$ gets smaller, more attention is put on heavy workload, during the process of training. According to our experimental results, setting $\gamma$ as the value indefinitely approaching zero can lead to the best accuracy. As analyzed before, when $\gamma$ indefinitely approaches zero, AHLF only optimizes the part with the highest prediction error. In this situation, PePNet also achieves high prediction accuracy without overall accuracy degradation, which verifies that AHLF’s use helps PePNet to avoid overly emphasizing sporadic patterns.
@@ III-E Ablation Experiment
We validate the effect of Periodicity-Perceived Mechanism and heavy-workload-focused loss function by comparing the performance of PePNet with those of PePNet- and PePNet. We show the performance of these models in Tab. II. Overall, Periodicity-Perceived Mechanism contributes more to the prediction accuracy than the AHLF. But the latter can be used to improve the accuracy of periodic and aperiodic time series, while Periodicity-Perceived Mechanism can only promote the accuracy of periodic and lax-periodic data. Furthermore, we test the prediction accuracy of PePNet on aperiodic data.
Performance of Periodicity-Perceived Mechanism. Dinda’s dataset and SMD dataset exhibit stricter periodicity than Alibaba2018 does [45]. Accordingly, the mechanism promotes the prediction accuracy most on Dinda’s dataset and SMD dataset. For datasets with lax periodicities, such as Alibaba2018-Memory, the mechanism also promotes overall and heavy-workload prediction accuracy. As for Alibaba2018-CPU, which has no significant periodicity on light load but has more significant periodicity on heavy workload, the mechanism can promote the heavy-workload prediction accuracy, while the performance of PePNet is close to the one of PePNet- on light workload. This observation verifies that the mechanism has little negative impact on aperiodic data.
Performance of AHLF. AHLF’s use can improve both overall and heavy-workload prediction accuracies on all three datasets, compared with PePNet. Except MAPE for heavy-workload prediction accuracy on memory usage of Alibaba2018, all of the evaluation metrics of PePNet are better than that of PePNet. As AHLF apparently reduces the error for extremely heavy workload in the long tail, which is also seen in Fig. 4(c), its use improves the accuracy for Dinda’s dataset and SMD dataset more that of Alibaba2018.
TABLE III: Compare the performance of PePNet on aperiodic and periodic data. The MSE are multiplied by 1000. The MAE are multiplied by 100.
@@
Performance of PePNet on aperiodic data. There is a major concern about whether PePNet still works well on aperiodic data. To test the prediction accuracy on aperiodic data, we divide Alibaba2018-CPU and Alibaba2018-Memory by periodicity and collect the prediction accuracy for periodic data and aperiodic data, respectively. In Tab. III, there is prediction accuracy for periodic data and aperiodic data. For convenience, the overall accuracy of both periodic data and aperiodic data is also listed at the bottom. On the whole, the accuracy for periodic data is slightly higher than the overall accuracy, while the accuracy for aperiodic data is slightly lower than the overall. However, there is a strange phenomenon MAPE of periodic data in Alibaba2018-Memory is bigger than the overall accuracy, while its MAE and MSE are much lower than the overall accuracy. That is because, in Alibaba2018-Memory dataset, the periodic workload is much smaller than the aperiodic one. Thus, even slight errors in periodic data prediction tend to become much larger after the division in the computation of MAPE.
@@ IV Related Work
Time series prediction aims to predict future time series based on the observed workload data. As an application branch of time series prediction, workload prediction aims to predict future workload based on observed ones.
Existing studies have studied a wide range of workload prediction tasks, including GPU workloads [19], CPU utilization [36], virtual desktop infrastructure pool workloads [41], disk states [11], and database query arrival rates [26]. These workload predictions are typically conducted to improve cloud system performance by optimizing the strategies of cloud resource allocation [7] and autoscaling [36].
In addition, several studies have proposed integrated frameworks that combine prediction models with other functional modules. For instance, Uncertainty-Aware Heuristic Search [24] integrates workload prediction with Bayesian optimization to explicitly model uncertainty and guide virtual machine provisioning strategies. SimpleTS [38] introduces an automated framework that selects appropriate prediction models for specific time series. However, these approaches primarily emphasize overall prediction accuracy, while largely overlooking accuracy under high-load conditions.
@@ V Conclusion
In this paper, we study how to improve overall workload prediction accuracy as well as heavy-workload one and propose PePNet. PePNet makes use of short-term dependent information, long-term tendency information and periodic information. Within PePNet, we propose two mechanisms for better prediction accuracy of workloads: 1) Periodicity-Perceived Mechanism to guide heavy-workload prediction, which can mine the periodic information and adaptively fuse periodic information for periodic, lax periodic and aperiodic time series; and 2) AHLF to offset the negative effect of data imbalance. We also provide theoretical support for the above design by: 1) providing a theoretically proven error bound of periodic information extracted by Periodicity-Perceived Mechanism; and 2) providing an automatic hyperparameter determination method for periodicity threshold $\mathcal{T}$. Compared with existing methods, extensive experiments conducted on Alibaba2018, Dinda’s dataset and SMD dataset demonstrate that PePNet improves MSE for overall workload prediction by 11.8% on average. Especially, PePNet improves MSE for heavy workload prediction by 21.0% on average.
Appendix A Proof of Theorem 1
According to the proposed periodicity-mining algorithm, when PePNet finds $k$, we have:
$\frac{E[a_{\tilde{t}}a_{t}]}{\sigma_{\tilde{t}}\sigma_{t}}>\mathcal{T}$ (13)
$E[a_{\tilde{t}}a_{t}]>\mathcal{T}*\sigma_{\tilde{t}}\sigma_{t}$ (14)
where $a_{i}=X_{i}-\mu_{i}$, $\tilde{t}=t-k$. $E[(X_{\tilde{t}}-X_{t})^{2}]$ can be transformed to:
Then, we substitute Eq. (14) into Eq. (15) to obtain:
Eq. (16) can be further reduced into:
Hence, the first part of Theorem 1 is proven.
When the time series is stationary, $\Delta\sigma^{2}$ and $\Delta\mu^{2}$ are zero. Then the upper bound can be reduced to $2(1-\mathcal{T})\sigma_{t-k}\sigma_{t}$.
Appendix B Proof of Theorem 2
The derivation of $l(Y,y)$ is:
Then, both numerator and dominator on the right-hand side are divided by $\exp(\frac{g(t^{*},\mathcal{P})}{\gamma})$ simultaneously to obtain:
$g(t^{*},\mathcal{P})$ is surely bigger than $g(t,\mathcal{P}),\forall t\neq t^{*}$, as the definition of $t^{*}$ suggests. Thus, $g(t,\mathcal{P})-g(t^{*},\mathcal{P})<0,\forall t\neq t^{*}$. When $\gamma(\gamma>0)$ infinitely approaches zero, $exp(\frac{g(t,\mathcal{P})-g(t^{*},\mathcal{P})}{\gamma})$ infinitely approaches zero as well. In this way, we have:
$\mathop{\lim}_{\gamma\rightarrow 0}\frac{\frac{\partial g(t^{*},\mathcal{P})}{\partial\mathcal{P}}+\sum_{t,t\neq t^{*}}^{T}[exp(\frac{g(t,\mathcal{P})-g(t^{*},\mathcal{P})}{\gamma})\frac{\partial g(t,\mathcal{P})}{\partial\mathcal{P}}]}{1+\sum_{t,t\neq t^{*}}^{T}exp(\frac{g(t,\mathcal{P})-g(t^{*},\mathcal{P})}{\gamma})}$ (20)
$=\frac{\partial g(t^{*},\mathcal{P})}{\partial\mathcal{P}}$